OpenStudy (anonymous):

Commutaive, Associative, Identity, Inverse. Name the property. y*1/y=1

OpenStudy (anonymous):

\[y \times \ \frac{ 1 }{ 5 }=1 \]

OpenStudy (anonymous):

\[y \times \frac{ 1 }{ y }=1\] correct one